OpenStudy (marissalovescats):

I need reminded of this if someone's willing. I know it's so simple, it's driving me crazy that I can't remember. Let u = <5, 6>, v = <-2, -6>. Find -2u + 5v

OpenStudy (jdoe0001):

\(\bf u = <5, 6>\ v = <-2, -6> \begin{cases} -2u\to -2<5,6>\to <-2\cdot 5,-2\cdot 6>\\ \bf 5v\to 5<-2,-6>\to <5\cdot -2,5\cdot -6> \end{cases} \\ \quad \\ {\color{brown}{ <a,b>+<c,d>\implies <a+c,b+d>}}\) meow!

OpenStudy (marissalovescats):

Ah right, thanks dear! What about this one? Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <7, 2>, v = <21, 6>

OpenStudy (marissalovescats):

Meow!

OpenStudy (johnweldon1993):

Well that would be our friend the dot product :D \[\large \vec u \cdot \vec v = <u_x,u_y> \cdot <v_x,v_y> = u_xv_x + u_yv_y\] if this comes out to 0, they are orthogonal

OpenStudy (johnweldon1993):

And to see if they are parallel we can use that same dot product \[\large u \cdot v = |u||v|cos\theta\] if they are parallel, then theta is 0 or 180...so we would just have \[\large u \cdot v = |u||v|\] or \[\large u \cdot v = -|u||v|\]

OpenStudy (marissalovescats):

Oh okay thanks!

OpenStudy (johnweldon1993):

And I feel as though "meow" is a reoccurring theme so "meow" welcome :D